A sodium nitrate solution has a molality of 0.50m. What is the mass percent of sodium nitrate?


1
Expert's answer
2021-04-08T12:11:29-0400

Molar mass of sodium nitrate is 84.9947 g


0.50×84.9947 =42.497 g


The total mass will be 42.497+103= 1042.497


Mass percent ="\frac{42.497}{1042.497}×100=4.08"%